Jane is 3 years older than jetty and 4 years younger than omoshi. How old is jetty if the sum of their ages is 70 years?
Answer by ikleyn(52798)

.
Jane is 3 years older than jetty and 4 years younger than omoshi.
How old is jetty if the sum of their ages is 70 years?
~~~~~~~~~~~~~~~~~~~~

Let Jane is x years old now.

Then Jetty is (x-3) years old, while Omoshi is (x+4) years old.


Write equation for the sum of ages

    x + (x-3) + (x+4) = 70.


Simplify and find x

    3x + 1 = 70

    3x = 70 - 1 = 69

     x          = 69/3 = 23.


Thus Jane is 23 years old.  Hence, Jetty is (23-3) = 20 years old.    ANSWER


CHECK.  Then Omoshi is (23+4) = 27 years old, and the sum of the three ages is

        20 + 23 + 27 = 70 years.  ! correct !

Solved.